OpenStudy (gavin39):

An airplane takes 4 hours to travel a distance of 2400 miles with the wind. The return trip takes 5 hours against the wind. Find the speed of the plane in still air and the speed of the wind.

OpenStudy (ovidia):

travel with the wind v+w=d/+1 v+w=3600/4 v+w=900 travel against the wind v-w=d/+2 v-w=3600/5 v-w=720 v+w=900 v-w=720 speed of the plane in the air is 810 mph speed of the wind is 90 mph